OpenStudy (dingdongwhothere):

question number 1Solve -2x = -19 * question number 2 Solve 9x = (-108) *

OpenStudy (dingdongwhothere):

@naate @AloneS @MysticFury @mathplane

OpenStudy (debbieg):

You need to put multiple questions in their own posts. For question number 1, do you know how to solve an equation where you have a coefficient in front of the variable? Divide both sides of the equation by the coefficient? A*x=C means that x=C/A Just apply that here.

OpenStudy (dingdongwhothere):

so just divide them @DebbieG

RhondaSommer (rhondasommer):

to solve for x, you need to divide on both sides of the equal sign. for ex) \[3x=9\]\[\frac{3x}{3}=\frac{9}{3}\]\[\frac{3x}{3}=1x~or~can~be~written~as~just~x\]\[\frac{9}{3}=3\]so x would equal 3 in this example.

RhondaSommer (rhondasommer):

in this equation (the first one), you would divide both sides by (-2)

OpenStudy (dingdongwhothere):

thannks